Adult tickets for a play cost $20 and child tickets cost $8. If there were 33 people at a performance and the theater collected $468 from ticket sales, how many adults and how many children attended the play?
Answer by checkley71(8403)

A+C=33 & 20A+8C=468
A=33-C NOW REPLACE "A" WITH (33-C) IN THE OTHER EQUATION & SOLVE FOR "C".
20(33-C)+8C=468
660-20C+8C=468
-12C=468-660
-12C=-192
C=-192/-12
C=16 ANSWER FOR THE NUMBER OF CHILDREN TICKETS SOLD.
A=33-16
A=17 ADULT TICKETS SOLD.
PROOF
20*17+8*16=468
340+128=468
468=468